Staff.

LBTO: Juan Carlos, Doug

Summary.

Details.

4:40PM power up WFS

Ocam frames ok.

ocam in warm mode.

Starting the night with MODS.

elevation 30.

Mirror position @30deg.

elevation30_mirror-pos.png

elevation30_act_25.png

ele82_mirror-pos.png

10:16 reconfigure LUCI..

05:24:36.298763| MAIN > Maximum mode index used to apply flat -> # 558....

ele90_mirror_pos.png

seeing 1.3"

Starting with 500 modes reconstractors.

10:34 osco sent.

rerotaotr shows to be at the limit.

reposition the rotator angle.

10:38 osco sent.

seeing 1" guider.

osco_20220122.png

wfs frames.

frames-20220122-bin1.png

10:52 ncpa on

SR 46%-40%

move rot to 60 deg.

data:

> Command line arguements processed
Set 0 Tracknum 20220122_054608 LuciName unknown
Set 1 Tracknum 20220122_054620 LuciName unknown
Set 2 Tracknum 20220122_054631 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_054608.log
Offset position:
telescope to Everyone (10:51 PM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_054938 LuciName unknown
Set 1 Tracknum 20220122_054950 LuciName unknown
Set 2 Tracknum 20220122_055002 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_054938.log
telescope to Everyone (10:54 PM)
1st position w/ NCPA:
telescope to Everyone (10:54 PM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_055303 LuciName unknown
Set 1 Tracknum 20220122_055315 LuciName unknown
Set 2 Tracknum 20220122_055327 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_055303.log
telescope to Everyone (10:57 PM)
Offset position w/ NCPA:
telescope to Everyone (10:57 PM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_055622 LuciName unknown
Set 1 Tracknum 20220122_055634 LuciName unknown
Set 2 Tracknum 20220122_055646 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_055622.log

11:04 adsec failure.

adsec_failure_20220122.png

adsc-stop/start

wfsarb stop/start.

11:12 osco sent

rot 60

ncpa 60nm astign

gopt gui off.

SR from modes ~0.7-0.6

no ncpa.

11:26 ncpa on

gopt gui on

start to skip frames....

loop opne.

goipt gui off ncpa off.

optimize gain + gopt.

seeing ~1.68"

ncpa on

skip-> loop open.

reclose

bin1_angle60_optimize.png

data..

st position, updated PA, w/o NCPA:
telescope to Everyone (11:21 PM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_062000 LuciName unknown
Set 1 Tracknum 20220122_062012 LuciName unknown
Set 2 Tracknum 20220122_062024 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_061959.log
telescope to Everyone (11:24 PM)
Offset pos., updated PA, w/o NCPA:
telescope to Everyone (11:24 PM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_062337 LuciName unknown
Set 1 Tracknum 20220122_062349 LuciName unknown
Set 2 Tracknum 20220122_062401 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_062337.log

change...recon to 400

link to new ACE table.

11:36pm preset osco

rot -40

gopt gui off.

new reco show a bite...

the reconstructor is not good !!!!!

stop AO...

change link to 500 modes.

11:50 PRESET IN BIN3

~11:57 TSS ENABLE/DISABLE JUST A KICK..? to check in the logs.

mag 16.3

seeing 1.3

optimize_bin3-20220122.png

clouds passing.

star off the pyramid.

cancel preset.

12:12 preset

mag 16.2

optimize gain,..

failed the second TT pass.

manual gains TT:4,ho1=1.2,ho2=0.15

seeing 1.2

luci stuck,... it seems that the process from the AO took too much time and luci hangs...

seeing 1.2 variable 1.4"

not double star.

a bit elongated, but this can come with a bad job of the gains in the TT

run manual gains

tt 10, ho14.2, ho2=0.25

round, not oscillation.....

12:44 bin3 fake with mag 14.51

seeing 1.4

mag 13.2

bin3_bright_.png

tn:20220122_075306.

1.----luci1.20220122.0082.

luci1.0090 with ncpa

Sr from modes ~0.6

Tn:080710

DATA

bin3 telemetry:
telescope to Everyone (12:43 AM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_074153 LuciName unknown
Set 1 Tracknum 20220122_074208 LuciName unknown
Set 2 Tracknum 20220122_074223 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_074153.log
telescope to Everyone (12:57 AM)
bin3 telemetry at G116-60:
telescope to Everyone (12:57 AM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_075535 LuciName unknown
Set 1 Tracknum 20220122_075544 LuciName unknown
Set 2 Tracknum 20220122_075553 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_075535.log

WITH NCPA

sr from modes ~0.4

bin3 telemetry w/ NCPA at G116-60:
telescope to Everyone (1:10 AM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_080929 LuciName unknown
Set 1 Tracknum 20220122_080937 LuciName unknown
Set 2 Tracknum 20220122_080945 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_080928.log

seeing 1.9" variable 1.6"

sr from modes 0.15

1:22am preset same target but with the Bin2.

mag 13.5

BIN2 826HZ

6 PHOT/SUBAP

THE OPTIMIZE GAIN IS SCANNING TOO LARGE VALUES

SEEING 1.3"

bin2_optimze_2ndpass.png

sr from modes 0.3"

done from flao and luci didn;t notice.

tn:20220122_083610

abort luci execution panel.

1:37 seeing 1.85

tn:_083757

seeing 1.5"

bin2_modalplot_01_41.png

tn:_084328

data

in2 telemetry w/o NCPA at G116-60:
telescope to Everyone (1:41 AM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_083906 LuciName unknown
Set 1 Tracknum 20220122_083917 LuciName unknown
Set 2 Tracknum 20220122_083926 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_083906.log

bin2 telemetry w/ NCPA at G116-60:
telescope to Everyone (1:48 AM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_084634 LuciName unknown
Set 1 Tracknum 20220122_084644 LuciName unknown
Set 2 Tracknum 20220122_084653 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_084634.log

1:44 ncpa on.

1:47 preset same target that fake bin3@380Hz.

1:55 preset. fake the magnitude.

mag 16.51

1:57 abort target, it is going to point toward the wind.

bin4

mag 16.5

seeing 1.18 variable 1.8"

3.5ph/subap

bin4_optimize.png

2:16

1.62"

bin4 telemetry at SBSS 1148_564:
telescope to Everyone (2:13 AM)
[telescope@obs1 ~]$ ./get_ao_telemetry

=> Command line arguements processed
Set 0 Tracknum 20220122_091144 LuciName unknown
Set 1 Tracknum 20220122_091155 LuciName unknown
Set 2 Tracknum 20220122_091207 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220122_091144.log

olmodes.

bin4_plotjitter_20220122_091144.png

residualmodes.

SR attenuation in H band due to TT jitter 0.012066733

bin4-plotjitter_residualmodes.png

2:20 preset

mag 13 fake for bin4

mag 14.9 13.7 pho/subap

tss enable/tss_disabled

2:29

optimize

0.12,0.10,0.01

bin4_optimizegain_secondstar.png

single star

_093733,_094142

sr from modes ~0.18"

2:43 preset.

mag 14.7 5pho/subap

bin3_pass3_2_49.png

_095046,_095414

3:06 frosty nebula.

rotator shows in the limit...

change rot.

hard to grab the nebula.

Rejected.

clouds coming.

3:19 authorize bincular luci.

preset ESM bin1.

stop because facing wind.

change to M51

3:28

seeing 1"

rerotator out of limit,,,, and went green again..

3:41 stop

Elliot access to telescope for refill.....

4:00am

clouds_20220122.png

4:10

restart M51

mag 17.3

seeing 1.78"

mag 16

optimize ok

esm_bin1_optimize.png

tn:_12225,_112439

luci_image_20220122.png

seeing 1.4"

bino mode.

binomode_20220122.png

TN:_114153

seeing 1.5"

some high wind speed above limit...

less cloud cover...

tn:_114951

4:54

seeing 1.3"

more clouds coming....

getting some extinsion..

5:10 loop open...

see pupils, but the loop doen;t hold. it seems background illumunation....

clouds_20220122_0525.png

5:31

reclose

tn:_123751

seeing guider dx 0.8"

getting corrections..

ocamviewer_0541.png

tn:_124731..loop pause.

done

5:57 preset

bin2.

1.5pho/subap oscillating...

skip frames....

optimize_gain_20220122_.png

too much extinsion....

waiting

call for the night

6:19am

-- JuanCarlosGuerra - 22 Jan 2022
I Attachment Action Size Date Who Comment
adsec_failure_20220122.pngpng adsec_failure_20220122.png manage 379 K 22 Jan 2022 - 06:05 JuanCarlosGuerra  
bin1_angle60_optimize.pngpng bin1_angle60_optimize.png manage 735 K 22 Jan 2022 - 06:37 JuanCarlosGuerra  
bin2_modalplot_01_41.pngpng bin2_modalplot_01_41.png manage 150 K 22 Jan 2022 - 08:41 JuanCarlosGuerra  
bin2_optimze_2ndpass.pngpng bin2_optimze_2ndpass.png manage 683 K 22 Jan 2022 - 08:34 JuanCarlosGuerra  
bin3_bright_.pngpng bin3_bright_.png manage 654 K 22 Jan 2022 - 07:52 JuanCarlosGuerra  
bin3_pass3_2_49.pngpng bin3_pass3_2_49.png manage 557 K 22 Jan 2022 - 09:50 JuanCarlosGuerra  
bin4-plotjitter_residualmodes.pngpng bin4-plotjitter_residualmodes.png manage 67 K 22 Jan 2022 - 09:27 JuanCarlosGuerra  
bin4_optimize.pngpng bin4_optimize.png manage 705 K 22 Jan 2022 - 09:17 JuanCarlosGuerra  
bin4_optimizegain_secondstar.pngpng bin4_optimizegain_secondstar.png manage 376 K 22 Jan 2022 - 09:36 JuanCarlosGuerra  
bin4_plotjitter_20220122_091144.pngpng bin4_plotjitter_20220122_091144.png manage 62 K 22 Jan 2022 - 09:24 JuanCarlosGuerra  
binomode_20220122.pngpng binomode_20220122.png manage 1 MB 22 Jan 2022 - 11:38 JuanCarlosGuerra  
clouds_20220122.pngpng clouds_20220122.png manage 1 MB 22 Jan 2022 - 11:01 JuanCarlosGuerra  
clouds_20220122_0525.pngpng clouds_20220122_0525.png manage 296 K 22 Jan 2022 - 12:26 JuanCarlosGuerra  
ele82_mirror-pos.pngpng ele82_mirror-pos.png manage 151 K 22 Jan 2022 - 03:18 JuanCarlosGuerra  
ele90_mirror_pos.pngpng ele90_mirror_pos.png manage 152 K 22 Jan 2022 - 05:26 JuanCarlosGuerra  
elevation30_act_25.pngpng elevation30_act_25.png manage 143 K 22 Jan 2022 - 02:32 JuanCarlosGuerra  
elevation30_mirror-pos.pngpng elevation30_mirror-pos.png manage 147 K 22 Jan 2022 - 02:30 JuanCarlosGuerra  
esm_bin1_optimize.pngpng esm_bin1_optimize.png manage 654 K 22 Jan 2022 - 11:21 JuanCarlosGuerra  
frames-20220122-bin1.pngpng frames-20220122-bin1.png manage 72 K 22 Jan 2022 - 06:13 JuanCarlosGuerra  
luci_image_20220122.pngpng luci_image_20220122.png manage 783 K 22 Jan 2022 - 11:26 JuanCarlosGuerra  
ocamviewer_0541.pngpng ocamviewer_0541.png manage 51 K 22 Jan 2022 - 12:42 JuanCarlosGuerra  
optimize_bin3-20220122.pngpng optimize_bin3-20220122.png manage 235 K 22 Jan 2022 - 07:03 JuanCarlosGuerra  
optimize_gain_20220122_.pngpng optimize_gain_20220122_.png manage 376 K 22 Jan 2022 - 13:11 JuanCarlosGuerra  
osco_20220122.pngpng osco_20220122.png manage 395 K 22 Jan 2022 - 05:47 JuanCarlosGuerra  
Topic revision: r2 - 22 Jan 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback